OpenStudy (anonymous):

Peter (P) and Jamie (j) have computer factories that are 132 miles apart. They both ship their completed computer parts to Diane(d) . Diane is 72 miles from peter and 84 miles from jamie. using points d,j,and p to form a triangle, fine pdj to the nearest ten minutes or nearest tenth of a degree. what is cos132?

OpenStudy (anonymous):

-.669

OpenStudy (amistre64):

-.669 dosnt look to be a correct answer :)

OpenStudy (amistre64):

maybe for cos132, but that seems spurious at best

OpenStudy (anonymous):

Person ask what cos 132 is, so I answer cos 132

OpenStudy (amistre64):

yeah, i see that :) but i think the cos132 might be wrong for the solution...gonna need to verify it

OpenStudy (amistre64):

we can modify the law of cosines to find the angle they are looking for

OpenStudy (amistre64):

132^2 = 84^2 + 72^2 -2(84)(72) cos(t); solve for t

OpenStudy (amistre64):

132^2 - 84^2 -72^2 ------------------- = cos(t) -2(84)(72)

OpenStudy (amistre64):

the angle asked for is the cos^-1 of that

OpenStudy (anonymous):

THAT SO COOL! :D What you did with division sign! :D I no think taht way! :D

OpenStudy (amistre64):

(132^2 - 84^2 -72^2)/(-2(84)(72)) = -0.428571429 t = 115.3769.... :)

OpenStudy (amistre64):

i find that way is alot clearer and faster than fighting the editor